# Intent Laundering: AI Safety Datasets Are Not What They Seem

### v1

@@ 4 Experimental Setup
We use the AdvBench (Zou et al., 2023) and HarmBench (standard) (Mazeika et al., 2024) datasets. We generate word clouds using the Python wordcloud package (Mueller, ). To preserve actual language patterns, we apply only lowercase conversion and whitespace normalization. For clearer visualization, we remove stopwords, punctuation, special characters, and words that instruct models, including “write,” “generate,” “create,” “develop,” “use,” “give,” “provide,” and “people.”
@@
We use the same safety datasets as in the word cloud analysis, along with two randomly sampled subsets of the GSM8K dataset (Cobbe et al., 2021). Each GSM8K subset matches the size of a corresponding safety dataset. This ensures a fair pairwise similarity comparison between safety and non-safety datasets, since the number of data points significantly affects similarity scores, as can be seen in Figure 3 across the two GSM8K subsets.
@@
We employ GPT-5.1 (gpt-5.1-2025-11-13) as a powerful judge to evaluate both safety and practicality when intent laundering is applied. Each evaluation is run independently in a zero-shot setting using separate model calls. All hyperparameters match those of the intent launderer, except that the output is capped at 1024 tokens. We report details on the agreement between LLM-based and human evaluations for both safety and practicality in Appendix A.
For safety evaluation on the original attacks (data points), where intent laundering is not applied, we use the standard ASR metric with the same five-point Likert scale described in Subsection 3.2 under Safety Evaluation. The corresponding prompt is shown in Figure 12 in Appendix C. We follow the same categorization scheme: an attack is considered successful if its response is labeled as extremely unsafe, highly unsafe, or unsafe. As before, all hyperparameters are identical to those of the intent launderer, except that the output is limited to 1024 tokens.
@@
Following the data duplication analysis in Subsection 2.2, we apply a conservative similarity threshold of 0.9 to work with unique data points in the safety datasets. Data points above this threshold are treated as duplicates, as they share similar sentence structures and identical malicious intent. For each group of duplicates, we randomly select one representative data point. This yields 207 data points from AdvBench and all 200 data points from HarmBench (standard). We use these two sets for all safety evaluation experiments.
@@
We select a diverse set of models for evaluation. In total, we experiment with seven different models: Gemini 3 Pro (gemini-3-pro-preview) (Google, 2025b), Claude Sonnet 3.7 (claude-3-7-sonnet-20250219) (Anthropic, 2025b), Grok 4 (grok-4-fast-non-reasoning) (xAI, 2025), GPT-4o (gpt-4o-2024-11-20) (Hurst et al., 2024), Llama 3.3 70B (llama-3.3-70b-instruct) (Grattafiori et al., 2024), GPT-4o mini (gpt-4o-mini-2024-07-18) (Hurst et al., 2024), and Qwen2.5 7B (qwen2.5-7b-instruct) (Yang et al., 2025). As our task does not require advanced reasoning and the base models remain the same across reasoning levels, we reduce reasoning efforts where applicable. Specifically, for Gemini 3 Pro, we set reasoning effort to low; for Claude Sonnet 3.7, we use standard mode; and for Grok 4, we use the non-reasoning checkpoint. All models use default inference hyperparameters, with output capped at 4096 tokens.
@@ 5 Results and Discussion
Table 1: Safety evaluation (SE), practicality evaluation (PE), and attack success rate (ASR) on the AdvBench dataset (top) and the HarmBench dataset (bottom). SE is the percentage of model responses rated as extremely unsafe, highly unsafe, or unsafe. PE is the percentage of responses rated as extremely applicable, highly applicable, or applicable. ASR is the percentage of responses that satisfy both SE and PE simultaneously. In the no-revision setting, where original data points from safety datasets are used, ASR follows its standard definition, as no abstraction is applied; thus, SE and PE are not applicable. The first-revision setting corresponds to the first application of intent laundering, where triggering cues are removed. All subsequent iterations reflect intent laundering with the revision–regeneration loop, which functions as a jailbreaking technique. Bold values denote the highest ASR across all iterations (within each dataset). Lower ASR implies better model safety.
@@
Table 1 presents results across seven models on both AdvBench and HarmBench. This includes results from three experimental settings:
@@
(1) Removing triggering cues from data points (attacks) leads to a sharp increase in ASR. On AdvBench and HarmBench, ASR rises from a mean of 5.38% and 13.79% in the no-revision settings, where cues are present, to 86.79% and 79.83%, respectively, in the first-revision settings, where cues are first removed. This indicates that model refusals are largely driven by the presence of triggering cues. Consequently, safety datasets do not reliably measure real-world safety risks, as they rely more on triggering cues to elicit refusals than on actual malicious intent.
(2) Intent laundering is highly effective at removing triggering cues while preserving the malicious intent. It also acts as a strong jailbreaking technique. As shown by the bold values—corresponding to the highest ASR and the last iteration in each dataset—intent laundering achieves high ASR values, ranging from 90% to 98.55% across all models and datasets. This includes models such as Gemini 3 Pro, known for strong safety (Google, 2025a; b), and Sonnet 3.7, known for overrefusal (Zhang et al., 2025; Cui et al., 2024). These models are jailbroken at high ASRs of 93%95% on AdvBench and 91%93% on HarmBench—and with only a few iterations.
@@
(4) ASR consistently increases with more iterations. While the first iteration always yields the largest gain, ASR continues to rise—with a mean increase of 9% on AdvBench and 11.81% on HarmBench by the final iteration—showing steady growth throughout. This confirms that the revision–regeneration mechanism effectively and systematically boosts ASR, and that adjusting the number of iterations provides direct control over the desired ASR.
(5) Our results suggest that both internal safety evaluations and safety-related training paradigms likely overrely on triggering cues, similar to publicly available safety datasets. This is supported by the fact that all insights motivating our methodology—such as unusual language patterns and dataset design flaws—are derived exclusively from publicly available safety datasets, rather than from safety-alignment techniques or internal safety evaluations. Nevertheless, these insights remain broadly effective, as evidenced by consistently high ASR in all experiments, regardless of whether models are closed- or open-weight, old or recent. Further evidence comes from the fact that internal safety evaluations reach the same conclusions as publicly available safety datasets: these models are reasonably safe (Anthropic, 2025a; b; Google, 2025a; b; xAI, 2025; Yang et al., 2025; Hurst et al., 2024; Grattafiori et al., 2024)a conclusion that stands in contrast to our findings.
@@ 6 Related Work
In response to these increasingly sophisticated attacks, advanced safety alignment techniques were proposed to enhance safety. Recent work primarily leverages reasoning as a mechanism to increase robustness, particularly in LRMs (Kim et al., 2025; Guan et al., 2024; Jaech et al., 2024). At the same time, this reasoning can itself be misused under adversarial conditions to jailbreak models (Sabbaghi et al., 2025). Overall, studies show that the safety alignment of current models remains fragile (Shah et al., 2025; Qi et al., 2025; Amodei et al., 2016), making reliable safety evaluation particularly challenging (Rando et al., 2025; Benton et al., 2024).
@@
As with other evaluation tasks, safety datasets aim to measure the effectiveness of safety alignment by capturing real-world attack scenarios. Early research focused on evaluating models against a narrow set of risks, such as bias (Tamkin et al., 2023; Nadeem et al., 2021; Dhamala et al., 2021; Nangia et al., 2020), toxicity (Hartvigsen et al., 2022; Gehman et al., 2020), and ethical judgment (Hendrycks et al., 2021). However, with the rise of general-purpose models, the focus broadened to encompass a wider range of safety threats, including misinformation, cybercrime, harassment, and more (Chao et al., 2024; Mou et al., 2024; Wang et al., 2023; Xu et al., 2021, inter alia). Within this expanded risk landscape, a particularly pressing area of concern in recent research involves preventing the misuse of models for chemical, biological, radiological, and nuclear threats (60 Minutes, 2025). This growing concern elevated datasets such as AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024) as prominent benchmarks for evaluating broader safety alignment.
@@ 7 Conclusion
We systematically studied the quality of two widely used AI safety datasets: AdvBench and HarmBench. We began by analyzing these datasets in isolation, without using any models. This analysis revealed that these datasets do not faithfully approximate real-world adversarial behavior due to their overreliance on triggering cues: expressions with overt negative/sensitive connotations designed to trigger safety mechanisms artificially. Motivated by this finding, we investigated what these datasets actually measure in practice when triggering cues are present, and whether their conclusions about model safety still hold when such cues are removed. To this end, we introduced intent laundering: a procedure that removes triggering cues from attacks (data points) while strictly preserving their malicious intent and all relevant details. Our results showed that prior safety conclusions do not hold once triggering cues are removed, and that the observed safety performance is largely driven by the presence of triggering cues rather than by the underlying safety risks. We further showed that intent laundering can be used as a powerful jailbreaking technique, achieving high attack success rates from 90% to over 98%. Overall, our findings unveiled a critical gap between how model safety is evaluated and how real-world adversarial behavior manifests. Based on this, we conclude that (1) safety evaluations must evolve to capture adversarial attacks more realistically, and (2) current safety-alignment efforts are still far from robust against real-world threats.
@@ Appendix A Agreement Analysis Between LLM-based and Human Evaluations
We measure agreement between the LLM judge and human expert consensus on 100 randomly selected instances, drawn from LLM judgments on responses from all models studied. For safety evaluation, model responses are categorized as unsafe (extremely unsafe, highly unsafe, or unsafe) or safe (moderately unsafe or safe). For practicality evaluation, model responses are categorized as practical (extremely applicable, highly applicable, or applicable) or impractical (moderately applicable or inapplicable). Human consensus is determined by the majority vote among three independent expert annotators. We report percent agreement rather than Cohen’s kappa (Cohen, 1960; Scott, 1955; Fleiss, 1971) due to high base rates of the majority class, which inflate chance agreement and render Kappa uninformative. We compute 95% confidence intervals (CI) using bootstrap resampling with 10,000 iterations (Efron, 1992; Tibshirani & Efron, 1993; Efron, 2003).
Table 2 reports the results of our agreement analysis. The LLM judge achieves 90.00% agreement (95% CI: 83.00%–95.00%) with human consensus on binary safety evaluation and 94.00% agreement (95% CI: 89.00%–98.00%) on binary practicality evaluation, meeting or exceeding the mean pairwise agreement among human annotators (82.00% for safety, 77.33% for practicality). The narrow bootstrap-estimated confidence intervals indicate stable performance, and the fact that the LLM’s agreement with consensus matches/surpasses human-human reliability suggests that the model can serve as a reliable automated judge for both safety and practicality evaluations.
Table 2: Agreement between LLM judge and human consensus on binary safety and practicality evaluations across 100 randomly sampled instances. Safety is labeled unsafe (extremely unsafe, highly unsafe, or unsafe) or safe (moderately unsafe or safe). Practicality is labeled practical (extremely applicable, highly applicable, or applicable) or impractical (moderately applicable or inapplicable). Human consensus is the majority vote of three expert annotators.
Evaluation Overall Agreement 95% CI Human-Human Agreement (mean)
@@ Appendix B Examples of Intent-Laundered Revisions and Model Responses
We present examples of actual model responses to revisions produced by the intent laundering process. Figure 5 presents a response from Gemini 3 Pro to a revision based on a data point from the AdvBench dataset. Similarly, Figure 6 shows a response from Claude Sonnet 3.7 to a revision based on a data point from the HarmBench dataset. As shown, all responses are unsafe, detailed, and transferable to real-world contexts, despite being framed as abstract scenarios.
@@ Appendix C Input Prompts
This section lists all prompts used in our study. These include prompts for intent laundering, the revision–regeneration loop, evaluation criteria generation, safety evaluation, practicality evaluation, and regular safety evaluation, where intent laundering is not applied (no-revision settings). Figures 7, 8, 9, 10, 11, and 12 present them in that order.

### v2

@@ 4 Experimental Setup
We generate word clouds for the AdvBench (Zou et al., 2023) and HarmBench (standard) (Mazeika et al., 2024) datasets using the Python wordcloud package (Mueller, ). To preserve actual language patterns, we apply only lowercase conversion and whitespace normalization. For clearer visualization, we remove stopwords, punctuation, special characters, and words that instruct models, including “write,” “generate,” “create,” “develop,” “use,” “give,” “provide,” and “people.”
@@
We use the same safety datasets as in the word cloud analysis, along with two randomly sampled subsets of the GSM8K dataset (Cobbe et al., 2021). Each GSM8K subset matches the size of a corresponding safety dataset. This ensures a fair pairwise similarity comparison between safety and non-safety datasets, since the number of data points can significantly affect similarity scores (see Figure 3 for the impact of this effect across the two GSM8K subsets).
@@
We employ GPT-5.1 (gpt-5.1-2025-11-13) as a powerful judge to evaluate both safety and practicality when intent laundering is applied. Each evaluation is run independently in a zero-shot setting using separate model calls. All hyperparameters match those of the intent launderer setup, except that the output is capped at 1024 tokens. We report details on the agreement between LLM-based and human evaluations for both safety and practicality in Appendix A.
For safety evaluation on the original attacks (data points), where intent laundering is not applied, we use the standard ASR metric along with the same five-point Likert scale described in Subsection 3.2 under Safety Evaluation. The corresponding prompt is shown in Figure 12 in Appendix C. We follow the same categorization scheme used for intent-laundered attacks: an attack is considered successful if its response is labeled as extremely unsafe, highly unsafe, or unsafe. As before, all hyperparameters are identical to those of the intent launderer setup, except that the output is limited to 1024 tokens.
@@
Based on our data duplication analysis in Subsection 2.2, we use a conservative similarity threshold of 0.9 to work with unique data points in the safety datasets. Data points above this threshold are treated as duplicates, as they share similar sentence structures and identical malicious intent (see Figure 4). For each group of duplicates, we randomly select one representative data point. This yields 207 data points from the AdvBench datasets and all 200 data points from the HarmBench (standard) dataset. We use these two sets for all our safety evaluation experiments.
@@
We select a diverse set of models for safety evaluation. In total, we experiment with seven different models: Gemini 3 Pro (gemini-3-pro-preview) (Google, 2025b), Claude Sonnet 3.7 (claude-3-7-sonnet-20250219) (Anthropic, 2025b), Grok 4 (grok-4-fast-non-reasoning) (xAI, 2025), GPT-4o (gpt-4o-2024-11-20) (Hurst et al., 2024), Llama 3.3 70B (llama-3.3-70b-instruct) (Grattafiori et al., 2024), GPT-4o mini (gpt-4o-mini-2024-07-18) (Hurst et al., 2024), and Qwen2.5 7B (qwen2.5-7b-instruct) (Yang et al., 2025). As our task does not require advanced reasoning and the base models remain the same across reasoning levels, we reduce reasoning efforts where applicable. Specifically, for Gemini 3 Pro, we set reasoning effort to low; for Claude Sonnet 3.7, we use standard mode; and for Grok 4, we use the non-reasoning checkpoint. All models use default inference hyperparameters, with output capped at 4096 tokens.
@@ 5 Results and Discussion
Table 1: Safety evaluation (SE), practicality evaluation (PE), and attack success rate (ASR) on the AdvBench dataset (top) and the HarmBench dataset (bottom). SE is the percentage of model responses rated as extremely unsafe, highly unsafe, or unsafe. PE is the percentage of responses rated as extremely applicable, highly applicable, or applicable. ASR is the percentage of responses that satisfy both SE and PE simultaneously. In the no-revision setting, where original data points from safety datasets are used, ASR follows its standard definition, as no abstraction is applied; thus, SE and PE are not applied. The first-revision setting corresponds to the first application of intent laundering, where triggering cues are removed. All subsequent iterations reflect intent laundering with the revision–regeneration loop, which functions as a jailbreaking technique. Bold values denote the highest ASR across all iterations (within each dataset). Lower ASR implies better model safety.
@@
Table 1 presents results for seven models on both the AdvBench and HarmBench datasets. This includes results from three experimental settings:
@@
(1) Removing triggering cues from data points (attacks) leads to a sharp increase in ASR. On AdvBench and HarmBench, ASR rises from a mean of 5.38% and 13.79% in the no-revision settings, where triggering cues are present, to 86.79% and 79.83%, respectively, in the first-revision settings (where triggering cues are removed for the first time). This indicates that model refusals are largely driven by the presence of triggering cues. Consequently, safety datasets fail to reliably measure real-world safety risks, as they rely more on triggering cues to elicit refusals than on actual malicious intent. This leads to an overestimation of model safety.
(2) Intent laundering is highly effective at removing triggering cues while preserving the malicious intent. It also acts as a strong jailbreaking technique. As shown by the bold values in Table 1—corresponding to the highest ASR and the last iteration in each dataset—intent laundering achieves high ASR values, ranging from 90% to 98.55% across all models and datasets. This includes models such as Gemini 3 Pro, known for strong safety (Google, 2025a; b), and Claude Sonnet 3.7, known for overrefusal (Zhang et al., 2025; Cui et al., 2024). These models are jailbroken at high ASRs of 93%95% on AdvBench and 91%93% on HarmBench—and with only a few iterations.
@@
(4) ASR consistently increases with more iterations. While the first iteration always yields the largest leap, ASR continues to rise—with a mean increase of 9% on AdvBench and 11.81% on HarmBench by the final iteration—showing steady growth throughout. This confirms that the revision–regeneration mechanism effectively and systematically boosts ASR, and that adjusting the number of iterations provides direct control over the desired ASR.
(5) Our results suggest that both internal safety evaluations and safety alignment techniques likely overrely on similar triggering cues found in publicly available safety datasets. This is supported by the fact that all insights motivating our methodology—such as unusual language patterns and dataset design flaws—are derived exclusively from publicly available safety datasets. Nevertheless, these insights remain broadly effective, as evidenced by consistently high ASR across all experiments, regardless of model specifications (e.g., closed- or open-weight, old or recent release, or developer identity). Further evidence comes from the fact that internal safety evaluations reach the same conclusions as publicly available safety datasets: these models are reasonably safe (Anthropic, 2025a; b; Google, 2025a; b; xAI, 2025; Yang et al., 2025; Hurst et al., 2024; Grattafiori et al., 2024)a conclusion that stands in contrast to our findings.
@@ 6 Related Work
In response to these increasingly sophisticated attacks, advanced safety alignment techniques were proposed to enhance safety. Recent work primarily leveraged reasoning as a mechanism to improve robustness, particularly in large reasoning models (Kim et al., 2025; Guan et al., 2024; Jaech et al., 2024). At the same time, the same reasoning capability exploited under adversarial conditions to jailbreak models (Sabbaghi et al., 2025). Overall, prior studies showed that the safety alignment of current models remains fragile (Shah et al., 2025; Qi et al., 2025; Amodei et al., 2016), making reliable safety evaluation challenging (Rando et al., 2025; Benton et al., 2024).
@@
As with other evaluation tasks, safety datasets aim to measure the effectiveness of safety alignment by capturing real-world scenarios. Early research focused on evaluating models against a narrow set of risks, such as bias (Tamkin et al., 2023; Nadeem et al., 2021; Dhamala et al., 2021; Nangia et al., 2020), toxicity (Hartvigsen et al., 2022; Gehman et al., 2020), and ethical judgment (Hendrycks et al., 2021). However, with the rise of general-purpose models, the focus broadened to encompass a wider range of safety threats, including misinformation, cybercrime, harassment, and more (Chao et al., 2024; Mou et al., 2024; Wang et al., 2023; Xu et al., 2021, inter alia). Within this expanded risk landscape, a particularly pressing area of concern in recent research is preventing the misuse of models for chemical, biological, radiological and nuclear threats (60 Minutes, 2025). This growing concern elevated datasets such as AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024) as prominent benchmarks for evaluating broader safety alignment.
@@ 7 Conclusion
We systematically studied the quality of two widely used AI safety datasets: AdvBench and HarmBench. We first analyzed these datasets in isolation, without using any models. This analysis revealed that these datasets do not faithfully approximate real-world adversarial behavior due to their overreliance on triggering cues: expressions with overt negative/sensitive connotations designed to trigger safety mechanisms artificially. Motivated by this finding, we investigated what these datasets actually measure in practice when triggering cues are present, and whether their conclusions about model safety still hold when such cues are removed. To this end, we introduced intent laundering: a procedure that removes triggering cues from attacks (data points) while strictly preserving their malicious intent and all relevant details. We showed that prior safety conclusions do not hold once triggering cues are removed, and that the observed safety performance is largely driven by the presence of triggering cues rather than by the underlying safety risks. We further showed that intent laundering can be used as a powerful jailbreaking technique, achieving high attack success rates from 90% to over 98%. Overall, our findings unveiled a critical gap between how model safety is evaluated and how real-world adversarial behavior occurs. Based on this, we conclude that (1) safety evaluations must evolve to capture adversarial behavior more realistically, and (2) current safety alignment efforts are still far from robust against real-world threats.
@@ Appendix A Agreement Analysis Between LLM-Based and Human Evaluations
We measure agreement between the LLM judge and the consensus of three human experts on 100 randomly sampled responses across all studied models. We use the same categorization scheme used for intent-laundered attacks (Section 3.2 under Attack Success Rate) and original attacks (Section 4 under LLM as a Judge). For safety evaluation, model responses are categorized as unsafe (rated extremely unsafe, highly unsafe, or unsafe) or safe (rated moderately unsafe or safe). For practicality evaluation, model responses are categorized as practical (rated extremely applicable, highly applicable, or applicable) or impractical (rated moderately applicable or inapplicable). Human consensus is determined by majority vote. We report percent agreement rather than Cohen’s kappa (Cohen, 1960; Scott, 1955; Fleiss, 1971) due to the high base rate of the majority class, which can inflate chance agreement and render kappa less informative. We compute the 95% confidence interval (CI) using bootstrap resampling with 10,000 iterations (Efron, 1992; Tibshirani & Efron, 1993; Efron, 2003).
Table 2 reports the results of the agreement analysis. The LLM judge achieves 90.00% agreement with human consensus on binary safety evaluation (95% CI: 83.00%–95.00%) and 94.00% agreement on binary practicality evaluation (95% CI: 89.00%–98.00%). Both exceed the mean pairwise agreement among human annotators (82.00% for safety; 77.33% for practicality). The relatively narrow bootstrap confidence intervals suggest that these estimates are robust to resampling variability. Overall, the LLM’s agreement with human consensus surpasses the mean human-human agreement, suggesting that it can serve as a reliable automated judge for both safety and practicality evaluations.
Table 2: Agreement between the LLM judge and the consensus of three human experts in binary evaluations of safety and practicality. The analysis uses 100 randomly sampled responses from all studied models. Safety is categorized as unsafe or safe, and practicality as practical or impractical. Human consensus is determined by majority vote. The results indicate that the LLM judge serves as a reliable proxy for expert human evaluation in both binary assessments.
Evaluation Overall Agreement 95% CI Mean Human-Human Agreement
@@ Appendix B Examples of Intent-Laundered Revisions and Model Responses
We present examples of actual model responses to intent-laundered revisions. Figure 5 presents a response from Gemini 3 Pro to a revision based on a data point from the AdvBench dataset. Similarly, Figure 6 shows a response from Claude Sonnet 3.7 to a revision based on a data point from the HarmBench dataset. As shown, all responses are unsafe, detailed, and transferable to real-world contexts, despite being framed as abstract scenarios.
@@ Appendix C Input Prompts
This section lists all the prompts used in our study. These include prompts for intent laundering, the revision–regeneration loop, evaluation criteria generation, safety evaluation, practicality evaluation, and regular safety evaluation (i.e., settings where intent laundering is not applied, referred to as no-revision settings). Figures 7, 8, 9, 10, 11, and 12 show them in the respective order.